# Super long-range kinks

### v1

@@ Abstract
In this work we investigate the presence of scalar field models supporting kink solutions with logarithmic tails, which we call super long-range structures. We first consider models with a single real scalar field and associate the long-range profile to the orders of vanishing derivatives of the potential at its minima. We then present a model whose derivatives are null in all orders and obtain analytical solutions with logarithmic falloff. We also show that these solutions are stable under small fluctuations. Remarkably, by investigating the forces between super long-range structures, we show that the kink-antikink force is stronger than the kink-kink one. Next, we study two-field models in which the additional field is used to modify the kinetic term of the other. By using a first-order formalism based on the minimization of the energy, we explore the situation in which one of the fields can be obtained independently from the other. Within this framework, we unveil how to smoothly go from long- or short- to super long-range structures.
@@ I Introduction
The kink solutions of the sine-Gordon model have exponential tails, that is, they attain their boundary values very fast, so we shall call them short range. The asymptotic behavior of the solutions is very important in the study of collisions, as it is associated with the interactions between the structures. Therefore, studying solutions with distinct tails is of interest. In this direction, one may seek for field profiles with long range. To achieve this, one may consider a limit in the double sine-Gordon model long0 whose solutions engender power-law tails. Since these solutions extend farther than the usual sine-Gordon ones, their interactions are stronger, so they are also called highly interactive. Over the years, several works dealing with long-range structures have appeared in the literature long0new ; long1new ; long2new ; long3new ; long1 ; long2 ; long3 ; forcelong ; forcelong2 ; long4 ; long5 ; long6 ; long7 .
@@
In this manuscript, we investigate the presence of kink structures with logarithmic tails, whose falloff is even slower than the usual long range ones. We call them super long-range structures. In Sec. II, we investigate how these structures arise under the action of models with a single real scalar field. We associate the super long-range character to the behavior of the derivatives of the potential calculated in its minima. We introduce a model in which all orders of the derivatives of the potential vanish at the minimum connected by the tail of the kink and obtain an analytical solution with logarithmic tails. The energy density is calculated, the linear stability is also investigated and the forces between pairs formed by super long-range kink-kink and kink-antikink are obtained. In Sec. III, we consider the inclusion of an additional scalar field that modifies the kinetic term of the other field. We use the BPS formalism and explore the situation in which one of the fields can be calculated independently from the other. Within this framework, we show how to smoothly go from long- to super long-range or from short- to super long-range configurations. We end the investigation in Sec. IV, where we present our final remarks and perspectives for future research.
@@ II Single scalar field model
The exponential tail in the solutions of the sine-Gordon model appears due to the finite non-null character of the classical mass. Since we are interested in obtaining the super long-range structures, let us first review the known long-range solutions. They appear in potentials that support null classical mass. In the double sine-Gordon model long0 , there is a special case that leads to the potential
@@
By expanding it asymptotically, one has $\phi(x)\approx\pi/2-1/x$. Therefore, the tail is of a power-law type, with falloff slower than the exponential one. This is the reason to call these structures long range or highly interactive. Other solutions with power-law tails can be found the in the literature long2 ; long3 . The energy density (3) takes the form
@@
where $n$ is a real parameter which obeys $n>2$. Regardless the value of $n$, the above expression supports minima at the very same location of the usual sine-Gordon potential. However, the derivatives at these points now obey $d^{m}V/d\phi^{m}\big{|}_{\phi=v_{i}}=0$ for $m=0,\ldots,\lceil{n-1}\rceil$. Thus, $n$ controls the orders of derivatives which vanish at the minima. In this situation, the first-order equation (4) with upper sign reads
@@
The presence of the power-law term in the above expression makes this approximation in the energy density being valid for distances smaller than the one required by the solution (16), which has a pure-logarithmic asymptotic behavior. We remark that, even though this falloff of the energy density is slower than the usual power-law one, it is still faster than the one related to the vacuumless model investigated in Refs. vacuum1 ; vacuum2 . We note, however, that contrary to the vacuumless solution, which is unlimited and arise from potentials with runaway minima, the solution (15) arise for the potential (14), which is somewhat similar to the well-known sine-Gordon model, so the asymptotic values of the solution are finite.
@@ II.2 Interkink forces
Since we are interested in calculating the forces between super long-range solutions, we have two configurations possible: kink-antikink ($KA$) and kink-kink ($KK$). Let us consider that the structures are symmetrically located around the origin, each one at the distance $L$ from the point $x=0$, where $L$ is much larger than the width of the solutions. In this situation, the separation is $2L$. Supposing that the structures are initially ($t=0$) at rest, the force in the kink at the left side is
@@
For the sine-Gordon solutions (7), we get $F=\pm 8e^{-2L}$, being positive (attractive) for the pair $KA$ and negative (repulsive) for configuration $KK$ forcesine . Notice that the force decays with an exponential function of the separation. For the usual long-range solution (10), we get the attractive force $F=8L^{-4}$ for the $KA$ pair and the repulsive force $F=-2L^{-4}$ for the $KK$ configuration, similarly to the results found in forcelong . Therefore, the power-law behavior of the solution also appears in the force.
In the case of the super long-range solution (15), we have shown that tail is logarithmic, extending farther than the power-law and exponential ones. In Fig. 7, we display the $KA$ pair formed by solutions in the interval $[-\pi/2,\pi/2]$ e and the $KK$ configuration formed by a solution in $[-\pi/2,\pi/2]$ joint with a solution in $[\pi/2,3\pi/2]$. Notice that, even though $L=50$ seems to be reasonable for our approximation in the case of exponential and power-law tails, it is not sufficient for the super long-range solutions, because the tails are too far from the minima connected by the solutions. This shows that the falloff of the solution (15) is indeed very slow, so one must consider higher separations to use the approximated solutions (30) and (31). For sufficiently large $L$ and $a>0$, the $KA$ force is
@@ III Two-field model
one obtains the first-order equations
$\frac{d\phi}{d\xi}=\pm\frac{\partial W}{\partial\phi},\qquad\frac{d\chi}{dx}=\pm\frac{\partial W}{\partial\chi}.$ (37)
The equation with upper/lower sign represents the increasing/decreasing solutions; they are related by the change $x\to-x$, so we only use the equations with the positive sign, for simplicity. In the first-order equation for $\phi$, $\xi$ represents the geometric coordinate that feeds the field $\phi$, given by
@@
Solutions of (37) minimize the energy of the system, $E=\left|W(v_{+},w_{+})-W(v_{-},w_{-})\right|$, where $v_{\pm}=\phi(\pm\infty)$ and $w_{\pm}=\chi(\pm\infty)$. We work with
@@ IV Outlook
In this manuscript, we have investigated one- and two-field models which support topological structures with logarithmic falloff, which we call super long range. First, we have considered the model described by the action (1). We have studied the classical mass associated to the minima of the potential to show that finite non-null mass leads to exponential tail, illustrated by the sine-Gordon potential (6). The usual long-range profile, described by power-law tails, appears in the case of null classical mass; this is seen in the special case of the double sine-Gordon model with the potential (9). However, we went deeper and investigated what happens as higher and higher orders of the derivatives of the potential vanish at the minima. To investigate this, we have consider the sine-Gordon like potential (12). We have concluded that, the higher the order of derivatives vanishing at the minima, the farther the tails extend, with the exponent controlling the power-law falloff depending on this feature.
Even though the potential (12) allows us to control the orders of vanishing derivatives at the minima through the parameter $n$, one cannot determine the behavior in the limit $n\to\infty$. To investigate this issue, we have introduced the potential (14), which engenders all orders of the derivatives null at the minima. We have obtained an exact solution and showed that its asymptotic behavior is described by logarithmic falloff. In the energy density, the tails possess terms which mix logarithimic and power-law demeanor; even though it is faster than the usual long-range configurations, it is still slower than the so-called vacuumless structures. We have also investigated the linear stability of the super long-range solutions, which is described a Schrödinger-like equation with stability potential of a volcano profile. We have showed that the zero mode does not support nodes, so the model is stable under small fluctuations. The forces between kink-antikink and kink-kink configurations were also investigated: we have concluded that they are described by the inverse of products between logarithmic and power-law functions of the separation. Remarkably, the exponents in the power-law contribution are different for each case, so the force is stronger for the $KA$ pair.
@@
whose asymptotic behavior is given by $\phi(x)\approx\pi/2-1/\ln(2\ln(2ax))$. Perhaps, a path to generalize the above result is by using the deformation method deform ; deform2 ; ganideform .
Moreover, in this work we have only investigated sine-Gordon-like potentials. Other perspectives include the study of polynomial potentials supporting super long-range configurations, both in one- and two-field models. Since solutions engendering logarithmic tails may exhibit forces with range and strength higher than usual, the study of their collisions is also of interest, as it may reveal distinct bounce windows.

### v2

@@ Abstract
In this work we investigate the presence of scalar field models supporting kink solutions with logarithmic tails, which we call super long-range structures. We first consider models with a single real scalar field and associate the long-range profile to the orders of vanishing derivatives of the potential at its minima. We then present a model whose derivatives are null in all orders and obtain analytical solutions with logarithmic falloff. We also show that these solutions are stable under small fluctuations. To investigate the forces between super long-range structures, we consider three methods and compare them. Next, we study two-field models in which the additional field is used to modify the kinetic term of the other. By using a first-order formalism based on the minimization of the energy, we explore the situation in which one of the fields can be obtained independently from the other. Within this framework, we unveil how to smoothly go from long- or short- to super long-range structures.
@@ I Introduction
The kink solutions of the sine-Gordon model have exponential tails, that is, they attain their boundary values very fast, so we shall call them short range. The asymptotic behavior of the solutions is very important in the study of collisions, as it is associated with the interactions between the structures. Therefore, studying solutions with distinct tails is of interest. In this direction, one may seek for field profiles with long range. To achieve this, one may consider the limit of the parameter of the double sine-Gordon model long0 which leads to a potential with null second derivative at the minima; in this case, the solutions engender the long-range character described by power-law tails. Since these solutions extend farther than the usual sine-Gordon ones, their interactions are stronger, so they are also called highly interactive. Over the years, several works dealing with long-range structures have appeared in the literature long0new ; long1new ; long2new ; long3new ; long1 ; long2 ; long3 ; forcelong ; forcelong2 ; long4 ; long5 ; long6 ; long67 ; long7 .
@@
In this manuscript, we investigate the presence of kink structures with logarithmic tails, whose falloff is even slower than the long-range ones with power-law asymptotic behavior. We call them super long-range structures. In Sec. II, we investigate how these structures arise under the action of models with a single real scalar field. We associate the super long-range character to the derivatives of the potential calculated in its minima. We introduce a model in which all orders of the derivatives of the potential vanish at the minimum connected by the tail of the kink and obtain an analytical solution with logarithmic tails. The energy density is calculated, the linear stability is also investigated and a discussion about the forces between pairs formed by super long-range kink-kink and kink-antikink is made. In Sec. III, we consider the inclusion of an additional scalar field that modifies the kinetic term of the other field. We use the BPS formalism and explore the situation in which one of the fields can be calculated independently from the other. Within this framework, we show how to smoothly go from long- to super long-range or from short- to super long-range configurations. We end the investigation in Sec. IV, where we present our final remarks and perspectives for future research.
@@ II Single scalar field model
The exponential tail in the solutions of the sine-Gordon model appears due to the finite non-null character of the classical mass. Since we are interested in obtaining the super long-range structures, let us first review the known long-range solutions. They appear in potentials that support null classical mass. In this direction, one may consider the double sine-Gordon model long0 , with $V(\phi)=(4/(1+4|\eta|))(\cos(\phi/2)-\eta\cos(\phi))$. One may take the special case $\eta=-1/4$ and perform the change $\phi\to 4\phi$ to get the potential $V(\phi)=4\cos^{4}(\phi)-3/2$. Inspired by this, we write
@@
By expanding it asymptotically, one has $\phi(x)\approx\pi/2-1/x$. Therefore, the tail is of a power-law type, with falloff slower than the exponential one. This is the reason to call these structures long range or highly interactive. Other solutions with power-law tails can be found in the literature long2 ; long3 ; long67 . The energy density (3) takes the form
@@
where $n$ is a real parameter which obeys $n>2$. Regardless the value of $n$, the above expression supports minima at the very same location of the usual sine-Gordon potential. However, the derivatives at these points now obey $d^{k}V/d\phi^{k}\big{|}_{\phi=v_{i}}=0$ for $k=0,\ldots,\lceil{n-1}\rceil$. Thus, $n$ controls the orders of derivatives which vanish at the minima. Notice that $k$ is integer, with maximum value determined by the ceiling function of argument $n-1$, which we have denoted by $\lceil{n-1}\rceil$. In this situation, the first-order equation (4) with upper sign reads
@@
The presence of the power-law term in the above expression makes this approximation in the energy density being valid for distances smaller than the one required by the solution (16), which has a pure-logarithmic asymptotic behavior. We remark that, even though this falloff of the energy density is slower than the usual power-law one, it is still faster than the one related to the vacuumless model, whose solutions are unlimited and arise from potentials with runaway minima, investigated in Refs. vacuum1 ; vacuum2 . However, we note that, contrary to the vacuumless configurations, the solution (15) emerges with the potential (14), which is somewhat similar to the well-known sine-Gordon model, so the asymptotic values of the solution are finite.
@@
The slow falloff of the super long-range configuration gives rise to a discussion about the thickness of the solution. As it is known in the literature, there are some ways to define thickness. One of them is using the energy density. Since the energy density vanishes asymptotically, we can define the thickness as $\Delta=2x_{t}$, where $x_{t}$ is the point in which the energy density obeys $\rho(x_{t})=\zeta$, with $\zeta$ being some value of interest near zero. Another way is by using some point $\tilde{x}$, such that $E_{\text{inside}}/E=\vartheta$, where $\vartheta$ represents the percentage of the total energy which is inside the interval $[-\tilde{x},\tilde{x}]$. In this situation, the thickness is $\Delta=2\tilde{x}$.
@@ II.2 Interkink forces
Since we are interested in calculating the forces between super long-range solutions, we have two configurations possible: kink-antikink ($KA$) and kink-kink ($KK$). Let us consider that the structures are symmetrically located around the origin, each one at the distance $L$ from the point $x=0$, where $L$ is much larger than the thickness of the solutions. In this situation, the separation is $2L$. Supposing that the structures are initially ($t=0$) at rest, the force in the kink at the left side is
@@
For the sine-Gordon solutions (7), we get $F=\pm 8e^{-2L}$, being positive (attractive) for the pair $KA$ and negative (repulsive) for configuration $KK$ forcesine . Notice that the force decays with an exponential function of the separation. For the specific long-range solution in (10), we get the attractive force $F=8L^{-4}$ for the $KA$ pair and the repulsive force $F=-2L^{-4}$ for the $KK$ configuration. Therefore, the power-law behavior of the solution also appears in the force. In Ref. forcelong , a comparison of these values with the ones obtained from accelerating kinks was made and, even though the dependence on the distance matches, the coefficient does not. To obtain a better coefficient, a method different from (30) and (31) was proposed. We shall refer to it as the gluing technique. It consists of cutting the left kink at $x=-L/2$ and the right kink at $x=L/2$ and gluing them with a function obtained from an approximation with adjusted coefficients of a Taylor expansion of (30) and (31) at the origin.
In the case of the super long-range solution (15), we have shown that tail is logarithmic, extending farther than the power-law and exponential ones. In Fig. 7, we display the $KA$ pair formed by solutions in the interval $[-\pi/2,\pi/2]$ and the $KK$ configuration formed by a solution in $[-\pi/2,\pi/2]$ joint with a solution in $[\pi/2,3\pi/2]$. Notice that, even though $L=50$ seems to be reasonable for our approximation in the case of exponential and power-law tails, it is not sufficient for the super long-range solutions, because the tails are too far from the minima connected by the solutions. This shows that the falloff of the solution (15) is indeed very slow, so one must consider higher separations to use the approximated solutions (30) and (31). For sufficiently large $L$ and $a>0$, the $KA$ force is
@@
To verify the expressions (32) and (33), we proceed similarly as in Ref. forcelong ; forcelong2 and compare them with the results obtained from accelerating kinks. To do so, we consider that the solutions used to form the configurations are well separated at time-dependent positions $-L(t)$ and $L(t)$. Following the aforementioned references, we model the accelerating kink at the left side by
$\phi(x,t)={\varphi}^{K}(x+L(t))$ (34)
for the same sectors of the potential studied with the approximation (30) and (31) plotted in Fig. 7. Notice that we have neglected the kink at the right side in the above approximation as we shall consider the interval $x\in[-L(t),0]$. By substituting the above field in the equation of motion (2), we get
$\frac{d^{2}{\varphi}^{K}}{dy^{2}}\pm A\frac{d{\varphi}^{K}}{dy}-\left.\frac{dV}{d\phi}\right|_{\phi={\varphi}^{K}}=0,$ (35)
in which $y=x+L(t)$. In this expression, we have disregarded the term proportional to $(dL/dt)^{2}$ and taken the acceleration as ±$\pm A=-d^{2}L/dt^{2}$. The positive/negative sign in (35) represents attractive/repulsive force associated with the $KA$/$KK$ configuration. Since the separation is large, we can assume that the acceleration is very small, $A\ll 1$. In this regime, we can use the BPS equation (4) to substitute $d{\varphi}^{K}/dy$ by $dW/d\phi|_{\phi={\varphi}^{K}}$ in the factor multiplying $A$ in (35), in which the auxiliary function $W(\phi)$ was introduced as usual in the BPS formalism bogo . From this, we can reduce (35) to the first order, in the form
$\frac{d{\varphi}^{K}}{dy}=\sqrt{2V({\varphi}^{K})\mp 2AW({\varphi}^{K})}.$ (36)
To keep the energy $E=\Delta W=M$, we fix the constant of the auxiliary function with $W(-\pi/2)=0$, so $W(\pi/2)=M$. By substituting the above expression into (29), one gets the force proportional to $MA$. Therefore, to obtain the force associated with accelerating kinks, we must integrate Eq. (36). For the super long-range solution (15), the potential is (14) and the tail can be obtained from (16). As it was done in Ref. forcelong , we take ${\varphi}^{K}$ and $\phi$ to diverge at the same points of the extrapolated tails, so we write
$\frac{\pi}{2}-{\varphi}^{K}\approx\frac{a}{\ln(2a(x+L))}.$ (37)
Since we are working within the region of the super long-range tail, we can take approximated forms for $V({\varphi}^{K})$ and $W({\varphi}^{K})$, so (36) becomes
$\frac{d{\varphi}^{K}}{dy}=\sqrt{4\left({{\varphi}^{K}}-\frac{\pi}{2}\right)^{4}e^{-\frac{2a}{|{{\varphi}^{K}}-\pi/2|}}\mp 2AM},$ (38)
where we have used the approximation $W({\varphi}^{K})=W(\pi/2)=M$ which, in the super long-range structure, is worse than in the cases of power-law and exponential tails. At this point we must be careful to properly choose the configuration before proceeding.
First, we consider the $KA$ pair. We then see that, at $x=-L$ ($y=0$), one has ${\varphi}^{K}=-\infty$ (using lateral limit in (37)) and, at $x=0$ ($y=L$), the derivative of ${\varphi}^{K}$ vanishes and the above expression leads us to $\left({\tilde{{\varphi}}^{K}}-\pi/2\right)^{4}e^{-\frac{2a}{|{\tilde{{\varphi}}^{K}}-\pi/2|}}=AM/2$. We perform the change of variables $\chi={\varphi}^{K}-\pi/2$ and integrate the above equation to get
$\int_{-\infty}^{\tilde{\chi}}\frac{d\chi}{\sqrt{4\chi^{4}e^{-\frac{2a}{|\chi|}}-2AM}}=L,$ (39)
where $\tilde{\chi}$ is the real solution of $\tilde{\chi}^{4}e^{-2a/|\tilde{\chi}|}=AM/2$ compatible with $A\ll 1$. The above integral is definite, so it leads to an expression in the form $f(A)=L$, from which we get $A=f^{-1}(L)$. Since the force between the $KA$ pair is attractive, we can write $F=Mf^{-1}(L)$. For the $KK$ pair, we can use similar arguments, with ${\varphi}^{K}=\pi/2$ at $x=0$, to obtain
$\int_{-\infty}^{0}\frac{d\chi}{\sqrt{4\chi^{4}e^{-\frac{2a}{|\chi|}}+2AM}}=L.$ (40)
In this situation, we can write the repulsive force as $F=-Mh^{-1}(L)$, where $A=h^{-1}(L)$ emerges from the above expression.
Both Eqs. (39) and (40) cannot be solved analytically. We then use numerical procedures to obtain $A$ and plot the force associated to the $KA$ and $KK$ pairs in the diamond-shaped points in Fig. 8. The curves of the accelerating-kink forces can be compared with the ones obtained with the approximation (30) and (31). Looking at the top panel of Fig. 8, we cannot see the curve representing the force (32) because it is too discrepant from values represented by the diamonds. This discrepancy can be seen in the inset.
Figure 8: The force associated to the kink-antikink ($KA$) pair (top panel) and the kink-kink ($KK$) pair (bottom panel) with respect to the parameter $L$, obtained from three different methods, for $a=1/2$. The solid lines in green stand for the forces (32) (top panel) and (33) (bottom panel), obtained with the approximation in Eqs. (30) and (31); there is no green line in the top panel because the force (32) is too discrepant from the ones obtained from the other methods. The inset shows the scale needed in the vertical axis to make the green line appear; it is a thousand times larger than the full panel. The diamonds in black stand for the forces obtained numerically from Eqs. (39) (top panel) and (40) (bottom panel) for $M=1.459$, which are fitted by the curves in red, described by $F^{KA}=0.33/(L^{2}\ln^{2}(L))$ and $F^{KK}=-60/(L^{2}\ln^{4}(20L))$. The lines in blue denote the forces in Eqs. (42) and (44), obtained from the gluing method. The scale of the vertical axis is represented by $\sigma^{KA}=3.0\times 10^{-9}$ in the top panel and $\sigma^{KK}=-1.5\times 10^{-7}$) in the bottom panel.
To find some other expression for the force that allows us to compare it with the values from the accelerating kinks or Eqs. (30) and (31), we use the gluing technique forcelong mentioned earlier which, in the region of interaction between the $KA$ pair, leads us to construct the field
$\phi^{KA}=\begin{cases}\frac{\pi}{2}-\frac{a}{\ln(2a|x+L|)},&x<-\frac{L}{2}\\ \frac{\pi}{2}-\frac{2a\ln(aL)-a}{2\ln^{2}(aL)}-\frac{2ax^{2}}{L^{2}\ln^{2}(aL)},&|x|\leq\frac{L}{2}\\ \frac{\pi}{2}-\frac{a}{\ln(2a|x-L|)},&x>\frac{L}{2}.\end{cases}$ (41)
In this case, the associated force (29) is
$F^{KA}=\frac{2a^{2}}{L^{2}\ln^{4}(aL)}.$ (42)
By comparing it with the expression in Eq. (32), we see that the coefficients and the exponent of the distance do not match. However, in the top panel of Fig. 8, we can see that the above expression approaches the curve obtained from the accelerating kinks better than (32), which does not appear in the plot because it is too discrepant. Moreover, by adjusting the parameters of the above expression, with $F^{KA}=0.33/(L^{2}\ln^{2}(L))$, we were able to get a good fit for the diamonds that describe the accelerating kinks.
For the $KK$ pair, the gluing method leads us to
$\phi^{KK}=\begin{cases}\frac{\pi}{2}-\frac{a}{\ln(2a|x+L|)},&x<-\frac{L}{2}\\ \frac{\pi}{2}+\frac{a(3\ln(aL)-1)x}{L\ln^{2}(aL)}-\frac{4a(\ln(aL)-1)x^{3}}{L^{3}\ln^{2}(aL)},&|x|\leq\frac{L}{2}\\ \frac{\pi}{2}+\frac{a}{\ln(2a|x-L|)},&x>\frac{L}{2},\end{cases}$ (43)
which, as for the $KA$, should only be used in the region of interaction. By using Eq. (29), we get the force
$F^{KK}=-\frac{9a^{2}}{2L^{2}\ln^{2}(aL)}.$ (44)
Interestingly, this expression only differs from (33) with respect to the coefficients; both of them appear in the bottom panel of Fig 8. By adjusting the coefficients of the above expression, taking $F^{KK}=-60/(L^{2}\ln^{4}(20L))$, one can fit the curve obtained by accelerating kinks.
We remark that there are other ways to construct the field profiles using the gluing method. For instance, one can consider the same expressions in Eq. (41) for $|x|>L/2$ and $\phi^{KA}(x)=\pi/2+\alpha+\beta x^{2}+\gamma x^{4}$ for $|x|\leq L/2$, with $\alpha=-a(4\ln^{2}(aL)-2\ln(aL)+1)/(4\ln^{3}(aL))$, $\beta=-2a(\ln(aL)-1)/(L^{2}\ln^{3}(aL))$ and $\gamma=-4a/(L^{4}\ln^{3}(aL))$ to keep the first and second derivatives continuous. The presence of the $x^{4}$ term which comes from the Taylor expansion around the origin does not modify the force (42). However, by considering extra terms for the $KK$ pair (43), i.e., $\phi^{KK}=\pi/2+\mu x+\lambda x^{3}+\sigma x^{5}$, with $\mu=-a(15\ln^{2}(aL)-6\ln(aL)+2)/(4L\ln^{3}(aL))$, $\lambda=-2a(5\ln^{2}(aL)-4\ln(aL)+2)/(L^{3}\ln^{3}(aL))$ and $\sigma=4a(3\ln^{2}(aL)-2\ln(aL)+2)/(L^{5}\ln^{3}(aL))$, for $|x|\leq L/2$, the force is not (44). Instead, we get $F^{KK}=-225a^{2}/(32L^{2}\ln^{2}(aL))$. To verify which approximation works better, one must numerically solve the time-dependent equation of motion (2). This is a deep question which we believe that has to be dealt in future works.
@@ III Two-field model
with the auxiliary function $W(\phi,\chi)$ being separable in $\phi$ and $\chi$, in the form $W(\phi,\chi)=G(\phi)+H(\chi)$, one obtains the first-order equations
$\frac{d\phi}{d\xi}=\pm\frac{dG}{d\phi},\qquad\frac{d\chi}{dx}=\pm\frac{dH}{d\chi}.$ (48)
The equations with upper/lower sign represent the increasing/decreasing solutions; they are related by the change $x\to-x$, so we only use the equations with the positive sign, for simplicity. In the first-order equation for $\phi$, $\xi$ represents the geometric coordinate that feeds the field $\phi$, given by
@@
In the previous works constr1 ; constr3 ; constr4 it was shown that $\xi$ may modify the kink significantly, changing its internal structure and/or its tail. We remark that, even though we are using a two-field model which is invariant under spatial translations, one may obtain a similar first-order equation by considering the presence of impurities, in which the aforementioned invariance is absent; see Ref. slawinska . In this situation, the coordinate $\xi$ can be related to the impurity. Solutions of (48) minimize the energy of the system, $E=\left|W(v_{+},w_{+})-W(v_{-},w_{-})\right|$, where $v_{\pm}=\phi(\pm\infty)$ and $w_{\pm}=\chi(\pm\infty)$. We work with
@@ IV Outlook
In this manuscript, we have investigated one- and two-field models which support topological structures with logarithmic falloff, which we call super long range. First, we have considered the model described by the action (1). We have studied the classical mass associated to the minima of the potential to show that finite non-null mass leads to exponential tail, illustrated by the sine-Gordon potential (6). The long-range profile, described by power-law tails, appears in the case of null classical mass; this is seen in the potential (9). However, we went deeper and investigated what happens as higher and higher orders of the derivatives of the potential vanish at the minima. To investigate this, we have consider the sine-Gordon-like potential (12). We have concluded that, the higher the order of derivatives vanishing at the minima, the farther the tails extend, with the exponent controlling the power-law falloff depending on this feature.
Even though the potential (12) allows us to control the orders of vanishing derivatives at the minima through the parameter $n$, one cannot determine the behavior in the limit $n\to\infty$. To investigate this issue, we have introduced the potential (14), which engenders all orders of the derivatives null at the minima. We have obtained an exact solution and showed that its asymptotic behavior is described by logarithmic falloff. In the energy density, the tails possess terms which mix logarithmic and power-law demeanor; even though it is faster than the usual long-range configurations, it is still slower than the so-called vacuumless structures. We have also investigated the linear stability of the super long-range solutions, which is described a Schrödinger-like equation with stability potential of a volcano profile. We have showed that the zero mode does not support nodes, so the model is stable under small fluctuations. The forces between kink-antikink and kink-kink configurations were also investigated using three different methods: we have concluded that they are described by the inverse of products between power-logarithmic and power-law functions of the separation. We have shown that the force (32) obtained with the ansatz (30) does not agree in functional dependence of the separation nor in coefficients when compared to the ones obtained from accelerating kinks and the gluing technique (see Fig. 8 and Eqs. (41) and (42)). The latter two seem to agree in functional dependence, only requiring adjustments in the coefficients, at least in the interval depicted in the top panel of Fig. 8. For the $KK$ pair, all the three methods lead to the same functional dependence, being distinguished only by the coefficients; see Eqs. (33) and (44). This was shown for the interval studied in Fig. 8. We remark that, even though we have found the forces using three methods, all of them involve approximations. Therefore, this issue brings to light an interesting perspective for future research, related to numerical simulation of dynamics using the time-dependent equation of motion.
@@
whose asymptotic behavior is given by $\phi(x)\approx\pi/2-1/\ln(2\ln(2ax))$. Perhaps, a path to generalize the above result is by using the deformation method deform ; deform2 ; ganideform1 ; ganideform2 . In this direction, the study of interkink forces may become even more trickier, as the multi-logarithmic tails fall off slower than the super long-range ones in (17).
Moreover, in this work we have only investigated sine-Gordon-like potentials. Other perspectives include the study of potentials involving polynomial functions, instead of sine-Gordon-like ones, supporting super long-range configurations, both in one- and two-field models. Since solutions engendering logarithmic tails may exhibit forces with range and strength higher than usual, the study of their collisions is also of interest, as it may reveal distinct bounce windows.